A car, travelling on a straight horizontal road, has 1.6 MJ of kinetic energy. It accelerates for 20 s until it has 2.5 MJ of kinetic energy. What is the average power output used to increase the kinetic energy of the car? *


1
Expert's answer
2021-01-16T17:21:28-0500

By the definition of the power, we have:


"P_{avg}=\dfrac{W}{t}=\dfrac{\Delta KE}{t},""P_{avg}=\dfrac{KE_f-KE_i}{t},""P_{avg}=\dfrac{2.5\cdot10^6\ J-1.6\cdot10^6\ J}{20\ s}=45000\ W=45\ kW."

Answer:

"P_{avg}=45000\ W=45\ kW."